# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

### v2

@@ 3.3 Theoretical Analysis of X-Boundary
In this subsection, we theoretically analyze the convergence rate of LLM from the perspective of the optimal transport theory (Solomon et al., 2020; Chuang et al., 2021; Weed & Bach, 2019). Specifically, we theoretically prove that X-boundary enables a faster learning speed of feature learning, which is verified in Fig. 4.
@@
Given a probability measure $\mu\in\textnormal{Prob}({\mathbb{R}}^{d})$ and a parameter $k\in{\mathbb{N}}$, the Wasserstein-$1$ $k$-variance is given as
$\mathrm{Var}_{k}(\mu)=\mathbb{E}_{S,\tilde{S}\sim\mu^{k}}\left[{\mathcal{W}}_{1}(\mu_{S},\mu_{\tilde{S}})\right],$ (6)
@@
$k$-variance measures structural properties of distribution beyond variance based on Wasserstein distances (Solomon et al., 2020). We theocratically analyze the learning trend of DNNs’ feature representations, which can be measured by the convergence rate of $k$-variance following (Weed & Bach, 2019; Solomon et al., 2020).
@@
(Proven in Appendix C) If $\phi_{\#}\mu$ is $(n,\Delta)$-clusterable, then for all $m\leq n(2\Delta)^{-2}$,
@@
Experimental Verification. Fig. 4 verifies that X-Boundary enables a faster learning speed of the training process. To this end, we fine-tune Llama-3-8B-Instruct and Qwen2.5-7B-Chat following settings in Section 2.1. Specifically, we set 0.1 and 0.55 of the training loss as thresholds to judge whether the training process has converged for Llama-3-8B-Instruct and Qwen2.5-7B-Chat, respectively. Based on this, Fig. 4 indicates that the proposed X-Boundary accelerates the converging process of 26.47% and 18.29% on Llama-3-8B-Instruct and Qwen2.5-7B-Chat, respectively.
Figure 4: The training curves of X-Boundary and without X-Boundary on Llama-3-8B-Instruct and Qwen2.5-7B-Chat.
@@ 4.1 Experimental Settings
To ensure fairness in comparison and consistency in experimental settings, we also implement X-Boundary on Llama-3-8B-Instruct, Qwen2.5-7B-Chat, and Mistral-7B-Instruct-v0.2, and evaluate it using the benchmarks described in Section 2.1. Additionally, to assess the effectiveness of X-Boundary across different sizes of LLMs, we implement it on Qwen2.5-14B-Chat. To construct the Separate Set, we sample 500 boundary-safe prompts from OR-Bench-80K (Cui et al., 2024), which have been filtered to avoid data contamination with the test set of OR-Bench. Next, we use GPT-4o to generate safe and helpful responses for these prompts, thus we get boundary-safe QA pairs. The retain set consists of our collected boundary-safe QA pairs, UltraChat (Ding et al., 2023), and refusal data points generated by the trained LLMs themselves. The erase set includes the harmful QA pairs for single-turn defense used in Zou et al. (2024) and the harmful QA pairs for multi-turn defense described in Section 2.1. More implementation details of X-Boundary are listed in Appendix B.3.
Table 2: Comparison of existing defense methods and X-Boundary on Qwen2.5-14B-Chat.
Methods Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
Vanilla 71.50 63.50 36.00 4.00 10.00 1.33 4.00 80.06 82.49 79.88
SFT 52.00 10.00 16.00 43.60 51.33 31.33 62.67 79.58 82.18 81.71
DPO 54.50 45.00 32.00 6.40 14.00 2.67 8.67 78.58 83.32 81.10
CB 23.50 4.50 8.00 43.60 51.33 32.00 64.33 79.64 82.56 82.93
X-Boundary 25.00 5.00 12.00 5.20 13.67 4.00 8.33 79.52 82.18 81.10
@@ 4.2 Main Results
The explicit formulation for boundary contributes to the precise distinction between harmful and safe representations. To investigate the effect of the explicit formulation for distinction boundary, we visualize the representation distribution of X-Boundary and without X-Boundary. Fig. 5 shows that, without X-Boundary, the boundary-safe representations close to harmful representations are mistakenly regarded as harmful ones. This demonstrates that LLMs fail to learn a boundary that exactly distinguishes safe and harmful representations, which supports our motivation of explicitly formulating the distinction boundary. With X-Boundary, harmful representations and boundary-safe representations are clearly separated as shown in Fig. 5, verifying that the proposed explicit formulation contributes to establishing a precise distinction boundary. Please refer to Appendix A.6 and Appendix A.7 for more detailed visualization and analysis of the representation distribution.
@@
X-Boundary maintains the lowest over-refusal rate while achieving SOTA defense against multi-turn jailbreaks. With a precise distinction boundary, X-Boundary reduces the ASR of ActorAttack by more than 40% while maintaining the increase in over-refusal rate on OKTest within 5% across three LLMs, as shown in Table 1. Specifically, on Llama-3-8B-Instruct, CB and X-Boundary both achieve the lowest ASR against ActorAttack, but X-Boundary demonstrates an average over-refusal rate that is lower by 20.05%. Similarly, on Qwen2.5-7B-Chat, X-Boundary’s average over-refusal rate is 58.50% lower than GA, which achieves the lowest ASR against Crescendo.
@@
X-Boundary rarely declines general capability. Table 1 shows that the decline of general capabilities caused by X-Boundary is generally no more than 0.5% compared to vanilla models, across the domains of general knowledge, mathematical ability, and coding ability. In contrast to SFT, which causes a 7% reduction in coding ability for Mistral-7B-Instruct-v0.2, X-Boundary achieves a lower ASR without compromising coding capability.
@@
X-Boundary successfully strikes a balance between robustness and usability. As a supplement to Table 1, Fig. 6 intuitively illustrates the trade-off between the ASR against multi-turn jailbreaks and the over-refusal rate. As the training process advances, the ASR steadily decreases, whereas the over-refusal rate progressively increases. Considering the two metrics comprehensively, X-Boundary appears in the lower-left corner of Fig. 6, indicating that it achieves a better balance compared to the baseline methods. In the same way, Fig. 8 demonstrates that X-Boundary also achieves a win-win outcome with robust defense and strong general capability. For specific cases of the defense performance and usability preservation of X-Boundary, please refer to Appendix D.
@@
X-Boundary is effective across different sizes of LLMs. Table 2 shows that, on Qwen2.5-14B-Chat, X-Boundary relatively reduces the ASR of three multi-turn attacks by more than 60%, while keeping the increase in over-refusal rate within 5% compared to the vanilla model. While X-Boundary and CB achieve comparable ASR, the over-refusal rate of X-Boundary is approximately 40% lower than that of CB. Compared with the performance on Qwen2.5-7B-Chat, the performance of X-Boundary on Qwen2.5-14B-Chat has not decreased.
@@
Figure 6: The trade-off between ASR of multi-turn jailbreaks and over-refusal rate on Llama-3-8B-Instruct. The data points were collected by sampling and evaluating at every 100 training steps.
Table 3: Comparison of existing defense methods and X-Boundary on reasoning models.
Models Methods Single & Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
HarmBench ActorAttack RedQueen XSTest OKTest OR-Bench PHTest AIME GPQA LiveCode
DeepSeek- R1-Distill- Llama-8B Vanilla 55.42 83.50 58.50 0.40 1.67 1.00 4.67 50.00 50.00 40.00
SFT 16.25 44.50 28.50 58.40 34.67 32.67 65.00 40.00 44.95 35.10
DPO 17.50 67.00 36.50 0.40 1.67 1.33 7.33 50.00 46.97 42.40
CB 52.80 77.50 55.00 1.60 3.33 3.67 8.00 46.67 46.97 40.65
X-Boundary 13.33 34.00 27.00 2.40 3.00 7.33 12.00 50.00 50.00 39.86
DeepSeek- R1-Distill- Qwen-7B Vanilla 60.42 79.00 60.50 1.60 2.00 1.33 2.67 53.30 48.99 39.76
SFT 48.33 59.50 42.50 7.60 9.00 11.00 12.67 46.67 48.99 36.44
DPO 52.92 74.00 48.50 1.60 3.67 3.00 8.67 53.33 50.00 40.08
CB 57.92 54.00 55.50 5.33 4.00 5.67 9.00 46.67 46.97 39.33
X-Boundary 20.80 43.50 36.00 2.40 3.00 7.33 12.00 53.33 50.51 40.02
Table 4: Ablation study on Qwen2.5-7B-Chat. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
@@
4.3 Defense Performance on Reasoning Models
Recently, Jiang et al. (2025); Zhou et al. (2025) have highlighted significant safety risks in large reasoning models, such as DeepSeek-R1 (Guo et al., 2025). Notably, the reasoning outputs of R1 models pose even greater safety concerns than their final answers. In this subsection, we explore the implementation of X-Boundary to enhance the safety defenses of reasoning models. As shown in Table 3, X-Boundary effectively reduces the ASR of both single-turn and multi-turn attacks, while not causing severe over-refusal issues or a decline in reasoning ability in distilled reasoning models. Surprisingly, the CB appears to be less effective for reasoning models, possibly because of differences in their representation space compared to standard models. According to the theoretical insights in Section 3.3, X-Boundary reduces the difficulty of training and accelerates the training process. Therefore, X-Boundary has an advantage over CB. The effect of defense methods on the reasoning ability of R1 models is discussed in Appendix A.2.
@@ 4.4 Ablation Study
We conduct ablation studies on the impact of multi-turn defense data, boundary-safe data, and separate loss. The results are illustrated in Table 4. More results on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2 are shown in Appendix A.4.
@@
Multi-turn defense data contribute to the reduction of ASR but intensify the over-refusal problem. With the multi-turn defense data described in Section 2.1 added into erase set, the ASR of ActorAttack is reduced from 63.00% to 15.50% on Qwen2.5-7B-Chat. However, the over-refusal rates in OR-Bench and PHTest increase about 30.00%. This highlights that over-refusal is a critical issue in multi-turn defense tasks, which cannot be overlooked and urgently requires resolution.
@@
Boundary-safe data can partially mitigate the over-refusal issue. Boundary-safe QA pairs added to the retain set significantly reduce the over-refusal rate on OR-Bench and PHTest but show limited effectiveness on XSTest and OKTest. This may be because the boundary-safe QA pairs are synthesized by LLMs, leading to effectiveness on OR-Bench and PHTest, which also use synthetic data for testing. In contrast, the test queries in XSTest and OKTest are manually crafted and may differ in distribution from the synthetic data, making it difficult to achieve effective generalization.
@@
Simply adjusting the size of boundary-safe data can not effectively balance ASR and over-refusal rate. Increasing the size of boundary-safe data can reduce the over-refusal rate, but it also leads to a sharp increase in ASR against multi-turn jailbreaks. Please see Appendix A.5 for more detailed results.
@@
Separate loss can further reduce the over-refusal rate. Unlike simply adding boundary-safe data, separate loss markedly reduces the over-refusal rate on both manually crafted and synthetically constructed benchmarks. Since the boundary-safe data shares the same source as OR-Bench, simply adding data is sufficient to reduce the over-refusal rate to a very low level, leaving little room for separate loss to make a noticeable impact. However, in the other three benchmarks, separate loss further reduces the over-refusal rate by an average of 9.75% compared to adding boundary-safe data.
5 Related Work
Multi-turn attack.
Several studies have explored the safety risks in multi-turn dialogue scenarios (Wang et al., 2025; Tong et al., 2024). For instance, Li et al. (2024a) employs human red teamers to uncover vulnerabilities in LLMs when subjected to multi-turn attacks. Jiang et al. (2024) crafts 40 multi-turn scenarios in which malicious intent is concealed under the guise of preventing harm. Yu et al. (2024), Zhou et al. (2024b) and Liu et al. (2024b) generate multi-turn jailbreak queries by breaking down the original malicious query into multiple less harmful sub-questions. Ren et al. (2024b) captures multi-turn attack clues by modeling a network of semantically linked actors. Yang et al. (2024b) and Russinovich et al. (2024) dynamically adjust the attack query based on the contextual feedback from victim LLMs, gradually steering benign initial queries toward more harmful topics throughout the conversation. In this paper, we evaluate the defense robustness of existing methods and X-Boundary against three types of multi-turn jailbreak attacks: ActorAttack (Ren et al., 2024b), RedQueen (Jiang et al., 2024), Crescendo (Russinovich et al., 2024).
Defenses for LLMs. Although defense methods for multi-turn jailbreak attacks are less explored in the literature, some existing approaches have proven effective against various single-turn attacks and have the potential to be adapted for multi-turn scenarios. These defense methods can be classified into the following categories: training LLMs to refuse harmful queries (Bai et al., 2022; Rafailov et al., 2024; Ouyang et al., 2022a; Yuan et al., 2024), training LLMs to prioritize safe instructions (Lu et al., 2024b; Wallace et al., 2024; Zhang et al., 2023), unlearning and editing harmful knowledge  (Lu et al., 2024a; Zhang et al., 2024b; Ren et al., 2024a; Qian et al., 2024a), prompt engineering (Xie et al., 2023; Zheng et al., 2024), and implementing input and output guardrails (Inan et al., 2023; Dubey et al., 2024) such as jailbreak detection (Hu et al., 2024a; Jain et al., 2023) input perturbation (Cao et al., 2023; Robey et al., 2023; Liu et al., 2024c). Several studies  (Li et al., 2024b; Zou et al., 2024, 2023; Qian et al., 2024b; Zhang et al., 2024a) also propose defense methods from the perspective of representation engineering, inspiring us to optimize LLMs in the representation space to strike a balance between defense robustness and LLM usability.
Decline in usability caused by defense methods. We assess the impact of defense methods on usability from two aspects: general capability degradation and over-refusal. General capability degradation, commonly known as the “alignment tax” (Ouyang et al., 2022b) phenomenon, has garnered widespread attention and has been extensively discussed in technical reports on LLMs (Dubey et al., 2024; Inan et al., 2023; Ren et al., 2024b; Li et al., 2024b; Hu et al., 2024b). Over-refusal refers to the unreasonable rejection of safe queries by LLMs (Varshney et al., 2023; Zhao et al., 2024; Zou et al., 2023; Arditi et al., 2024; Cao et al., 2024). Bianchi et al. (2023) discover that excessive safety-tuning makes LLMs refuse entirely safe prompts if they superficially resemble unsafe ones. Röttger et al. (2023), Shi et al. (2024), Cui et al. (2024), and An et al. (2024) employ linguistic techniques or automatic pipelines to generate seemingly unsafe prompts for evaluating LLMs’ over-refusal behavior.

### v3

@@ 3.3 Theoretical Analysis of X-Boundary
In this subsection, we theoretically analyze the convergence rate of LLM from the perspective of the optimal transport theory solomon2020k; chuang2021measuring; weed2017sharp. Specifically, we theoretically prove that X-boundary enables a faster learning speed of feature learning, which is verified in Fig. 5.
@@
Given a probability measure $\mu\in\textnormal{Prob}({\mathbb{R}}^{d})$ and a parameter $k\in{\mathbb{N}}$, the Wasserstein-$1$ $k$-variance is
@@
where $\mu_{S}=\frac{1}{k}\sum_{i=1}^{k}\delta_{x_{i}}$ for $x_{i}\overset{\textnormal{i.i.d.}}{\sim}\mu$.
$k$-variance measures structural properties of distribution beyond variance based on Wasserstein distances solomon2020k. We theoretically analyze the learning trend of DNN feature representations, which can be measured by the convergence rate of $k$-variance following weed2017sharp; solomon2020k.
@@
(Proven in Appendix E) If $\phi_{\#}\mu$ is $(n,\Delta)$-clusterable, then for all $m\leq n(2\Delta)^{-2}$,
@@
Experimental Verification. Fig. 5 verifies that X-Boundary enables a faster learning speed of the training process. To this end, we fine-tune Llama-3-8B-Instruct and Qwen2.5-7B-Instruct following the settings in Section 2. Specifically, we set 0.1 and 0.55 of the training loss as thresholds to judge whether the training process has converged for Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, respectively. Based on this, Fig. 5 indicates that the proposed X-Boundary accelerates the converging process of 26.47% and 18.29% on Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, respectively.
Figure 5: The training curves of X-Boundary and without X-Boundary on Llama-3-8B-Instruct and Qwen2.5-7B-Instruct.
@@ 4 Experiments
Methods Single-Turn ASR (%) $\downarrow$ Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
  GCG PAIR PAP ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Llama-3-8B-Instruct
Vanilla 31.00 18.00 15.00 58.50 25.00 34.00 6.80 9.00 8.00 13.67 68.30 79.08 59.18
SFT 6.50 13.50 1.50 19.50 0.50 8.00 27.20 42.33 22.00 57.33 68.17 76.19 54.27
DPO 8.50 11.00 3.00 17.50 5.00 14.00 20.00 28.33 17.33 41.00 68.01 75.59 58.54
GA 18.00 11.50 3.50 38.50 1.50 12.00 10.80 15.00 13.33 35.33 68.25 77.86 62.20
CB 2.00 12.00 1.00 16.50 0.50 10.00 23.60 27.67 36.00 52.00 67.66 78.47 59.76
X-Boundary 1.50 10.00 1.00 16.50 1.00 10.00 8.40 14.00 8.00 28.67 67.94 78.70 59.76
Qwen2.5-7B-Instruct
Vanilla 76.00 48.50 51.50 76.00 39.50 62.00 6.00 19.33 1.67 5.67 74.26 80.67 81.71
SFT 48.50 39.50 15.50 21.00 6.00 18.00 46.00 57.67 29.33 53.67 74.30 76.42 77.44
DPO 46.50 48.00 21.50 38.00 12.00 24.00 21.60 25.67 11.67 32.33 73.63 80.97 80.49
GA 54.00 35.00 9.50 38.00 21.00 12.00 58.33 70.00 67.67 85.33 74.58 80.43 79.27
CB 22.00 27.50 10.50 15.50 5.50 12.00 20.60 26.00 34.00 43.67 74.21 80.36 81.10
X-Boundary 23.00 26.00 8.50 17.50 7.50 16.00 10.40 16.67 5.33 15.00 74.17 80.52 81.10
Mistral-7B-Instruct-v0.2
Vanilla 83.50 60.50 61.00 70.00 49.50 40.00 10.00 21.00 4.33 13.00 59.98 45.34 34.76
SFT 38.50 48.00 34.00 37.50 22.00 18.00 53.60 42.00 29.33 58.67 58.94 41.55 27.44
DPO 36.00 47.00 42.50 44.50 19.00 28.00 25.20 38.67 20.33 37.67 58.79 43.21 34.76
GA 48.00 32.50 25.00 24.00 9.00 10.00 38.40 50.67 35.67 71.33 60.13 45.00 34.76
CB 31.00 36.50 30.50 15.00 11.50 12.00 45.20 39.33 55.00 50.00 59.91 46.63 33.54
X-Boundary 34.50 35.00 30.00 16.00 13.50 14.00 19.20 23.33 10.34 26.33 59.83 45.34 36.59
Table 1: Comparison of existing defense methods and X-Boundary.
@@ 4.1 Experimental Settings
To ensure fairness in comparison and consistency in experimental settings, we implement four baseline methods and X-Boundary on Llama-3-8B-Instruct, Qwen2.5-7B-Instruct, and Mistral-7B-Instruct-v0.2, and evaluate them using HarmBench dataset harmbench and the metrics described in Section 2. Additionally, to assess the effectiveness of X-Boundary across different sizes of LLMs, we implement it on Qwen2.5-14B-Instruct. To construct the Separate Set, we sample 500 boundary-safe prompts from OR-Bench-80K orbench, which have been filtered to avoid data contamination with the test set of OR-Bench. Next, we use GPT-4o to generate safe and helpful responses for these prompts, thus we get boundary-safe QA pairs. The retain set consists of boundary-safe QA pairs, UltraChat ultrachat, and refusal data points generated by the trained LLMs themselves. The erase set includes the harmful QA pairs for single-turn defense used in circuit_breaker and the harmful QA pairs for multi-turn defense described in Section 2. Evaluation and implementation details of X-Boundary are listed in Appendix D.5 and D.3, respectively.
@@ 4.2 Main Results
The explicit formulation for boundary contributes to the precise distinction between harmful and safe representations. To investigate the effect of the explicit formulation for distinction boundary, we visualize the representation distribution of X-Boundary and without X-Boundary. Fig. 6 shows that, without X-Boundary, the boundary-safe representations close to harmful representations are mistakenly regarded as harmful ones. This demonstrates that LLMs fail to learn a boundary that exactly distinguishes safe and harmful representations, which supports our motivation of explicitly formulating the distinction boundary. With X-Boundary, harmful representations and boundary-safe representations are clearly separated as shown in Fig. 6, verifying that the proposed explicit formulation contributes to establishing a precise distinction boundary. Please refer to Appendix C.10 and C.11 for more detailed visualization of the representation distribution.
@@
X-Boundary maintains the lowest ORR while achieving SOTA defense against both single-turn and multi-turn jailbreaks. With a precise distinction boundary, X-Boundary relatively reduces single-turn and multi-turn ASR by more than 40% while maintaining the increase in ORR on OKTest within 5% across three LLMs, as shown in Table 1. Specifically, on Llama-3-8B-Instruct, CB and X-Boundary both achieve the lowest ASR against ActorAttack, but X-Boundary demonstrates an average ORR that is lower by 20.05%. Similarly, on Qwen2.5-7B-Instruct, X-Boundary’s average ORR is 58.50% lower than GA, which achieves the lowest ASR against Crescendo.
@@
X-Boundary rarely declines general capability. Table 1 shows that the decline of general capabilities caused by X-Boundary is generally no more than 0.5% compared to vanilla models, across the domains of general knowledge, mathematical ability, and coding ability. In contrast to SFT, which causes a 7% reduction in coding ability for Mistral-7B-Instruct-v0.2, X-Boundary achieves a lower ASR without compromising coding capability. More evaluations of single-turn defense are listed in Appendix C.3.
@@
X-Boundary successfully alleviates the trade-off between robustness and usability. As a supplement to Table 1, Fig. 7 intuitively illustrates the trade-off between ASR against jailbreaks and ORR. Considering the two metrics comprehensively, X-Boundary appears in the lower-left corner of Fig. 7 and increases the hypervolume, i.e., the volume of the dominated space between the Pareto front and a predefined reference point, by 13.13% and 10.03% in OKTest and PHTest, respectively. The results indicate that X-Boundary significantly advances the Pareto frontier and mitigates the trade-off between ASR and ORR compared to the baseline methods. In the same way, Fig. 10 in Appendix C.5 demonstrates that X-Boundary also achieves a win-win outcome with robust defense and strong general capability. For specific cases of the defense performance and usability preservation of X-Boundary, please refer to Appendix F.
@@
X-Boundary is effective across different sizes of LLMs. Table 5 in Appendix C.2 shows that, on Qwen2.5-14B-Instruct, X-Boundary relatively reduces the ASR by more than 60%, while keeping the increase in ORR within 5% compared to the vanilla model. Although X-Boundary and CB achieve comparable ASRs, the ORR of X-Boundary is approximately 40% lower than that of CB. Compared with the performance on Qwen2.5-7B-Instruct, those of X-Boundary on Qwen2.5-14B-Instruct is stable and has not decreased.
@@
Figure 7: The trade-off between ASR of jailbreaks and ORR. The data points are collected by sampling and evaluating every 100 training steps.
Figure 8: Comparison of existing defense methods and X-Boundary on DeepSeek-R1-distilled models.
@@
Table 2: Ablation study on Qwen2.5-7B-Instruct. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
4.3 Performance on Large Reasoning Models
Recently, several studies safechain; R1_assessment have highlighted significant safety risks in the outputs of large reasoning models (LRMs), particularly during the thinking process. Enhancing the security of LRMs, such as DeepSeek-R1 guo2025deepseek, has become an urgent priority. In this section, we evaluate the performance of X-Boundary and baseline methods on two LRMs: DeepSeek-R1-distilled-LLaMA-8B and DeepSeek-R1-distilled-Qwen-7B. The evaluation of defense performance and over-refusal adopts the same datasets and metrics as Section 2 described. To assess general capability, we replace the previous datasets with more challenging benchmarks that test reasoning capability (RC), namely AIME2024, GPQA, and LiveCodeBench. The detailed evaluation settings and analysis of RC are listed in Appendix C.4.
As shown in Fig. 8, both CB and DPO exhibit marginal defense effectiveness on LRMs, reducing the average ASR by only around 10% on the Distilled-Qwen model. Although SFT still demonstrates robust defense on LRMs, it causes a degradation of over 5% in RC and leads to a significant increase in the average ORR. In contrast, X-Boundary achieves outstanding defense performance while maintaining the average ORR below 10% and preserving 99% RC. This result may be attributed to the theoretical analysis in Section 3.3, which suggests that X-Boundary reduces the difficulty of training and facilitates faster convergence within the complex representation space of LRMs.
@@ 4.4 Ablation Study
We conduct ablation studies on the impact of multi-turn defense data, boundary-safe data, and separate loss. The results are illustrated in Table 2. Ablation Studies on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2 are shown in Appendix C.8. Please see Appendix C.6 and C.7 for ablation studies on three terms of loss and sensitivity analysis on hyper-parameters $\alpha$ and $\beta$, respectively.
@@
Multi-turn defense data contribute to the reduction of ASR but intensify the over-refusal problem. With the multi-turn defense data described in Section 2 added into the erase set, the ASR of ActorAttack is reduced from 63.00% to 15.50% on Qwen2.5-7B-Instruct. However, the ORRs in OR-Bench and PHTest increase by about 30.00%.
@@
Boundary-safe data can partially mitigate the over-refusal issue. Boundary-safe QA pairs added to the retain set significantly reduce the ORR on OR-Bench and PHTest but show limited effectiveness on XSTest and OKTest. This may be because the boundary-safe QA pairs are synthesized by LLMs, leading to effectiveness on OR-Bench and PHTest, which also use synthetic data for testing. In contrast, the test queries in XSTest and OKTest are manually crafted and may differ in distribution from the synthetic data, making it difficult to achieve effective generalization.
@@
Simply adjusting the size of boundary-safe data can not effectively balance ASR and ORR. Increasing the size of boundary-safe data can reduce the ORR, but it also leads to a sharp increase in ASR against jailbreaks. Please see Appendix C.9 for more detailed results.
@@
Separate loss can further reduce the ORR. Unlike simply adding boundary-safe data, separate loss markedly reduces the ORR on both manually crafted and synthetically constructed benchmarks. Since the boundary-safe data shares the same source as OR-Bench, simply adding data is sufficient to reduce the ORR to a very low level, leaving little room for separate loss to make a noticeable impact. However, in the other three benchmarks, separate loss further reduces the ORR by an average of 9.75%.